Fast Inc. has no debt and a WACC of 9%. The company now changes its capital structure to a debt-to-equity ratio of 0.5. The interest rate on debt is 4.5%, the new cost of equity is 10.51% and the tax rate is 33%.

What is the company's new WACC?
Answer #1

Weight of equity= 1 / (1 + D/E)

Weight of equity = 1 / (1 + 0.5)

Weight of equity = 0.6667

Weight of debt = 1- 0.6667 = 0.3333

New WACC = Weights of debt*after tax cost of debt + weight of equity*cost of equity

New WACC = 0.3333*0.045*(1 - 0.33) + 0.6667*0.1051

New WACC = 0.010049 + 0.07007

New WACC = 0.0801 or 8.01%